# Learning A Physical-aware Diffusion Model Based on Transformer for Underwater Image Enhancement

### v1

@@ 3.2 Physics Prior Generation
$\hat{I}=GT\cdot\mathcal{T}(I)+(1-\mathcal{T}(I))\mathcal{B}(I),$ (9)
@@
We constrain the PPG branch by supervised manner. The physics reconstruction loss can be expressed as:
$\mathcal{L}_{p}=||\hat{I}-I||_{1},$ (10)
Finally, we use a Conv and GAP to extract physics prior information $\mathbf{T^{c}}$ and $\mathbf{B^{c}}$ to guide the diffusion process.
3.3 Physics-aware Diffusion Transformer
@@
Physics-aware diffusion transformer (PDT) branch aims to exploit the extracted physics prior information from PPG branch for better guiding underwater image restoration. Fig. 1 shows the overall of PDT branch, which is a U-shaped structure with physics-aware diffusion transformer blocks (PDTB). We also utilize skip connections to connect the features at the same level.
@@
Given a degraded underwater image as condition image $x_{c}\in{R}^{H\times W\times 3}$, we first obtain the noise sample $x_{t}\in{R}^{H\times W\times 3}$ at time step $t$ according to the forward process, and concatenate $x_{t}$ and $x_{c}$ at channel dimension to obtain the input of PDT branch. Then, we obtain the embedding features $\mathbf{F}\in{R}^{H\times W\times C}$ through convolution, where C means the number of channel. $\mathbf{F}$ is encoded and decoded by PDTB, which consists of physics-aware self-attention (PA-SA) and multi-scale dynamic feed-forward network (MS-FFN). In our PA-MSA, We first embed the time embedding $\mathbf{T}$ into the input features $\mathbf{F}$ :
$\tilde{\mathbf{F}}=Norm(\mathbf{F})+\mathbf{T},$ (11)
where $Norm$ means layer normalization. Then, we integrated $\mathbf{T^{c}}$ and $\mathbf{B^{c}}$ as the corresponding dynamic modulation parameters to generate transmission-aware feature $\mathbf{F_{t}}$ and light-aware feature $\mathbf{F_{b}}$:
$\mathbf{F_{t}}$ $=Li(\mathbf{T^{c}})\odot\tilde{\mathbf{F}},$ (12)
$\mathbf{F_{b}}$ $=Li(\mathbf{B^{c}})\odot\tilde{\mathbf{F}},$
where $Li()$ means linear layer. Afterward, we aggregate the obtained embeddings by projecting $\mathbf{F_{b}}$ into query $\mathbf{Q}=\mathbf{W}_{d}\mathbf{W}_{p}\mathbf{F}_{\mathbf{b}}$, $\mathbf{F_{t}}$ into key $\mathbf{K}=\mathbf{W}_{d}\mathbf{W}_{p}\mathbf{F}_{\mathbf{t}}$ and transforming $\tilde{\mathbf{F}}$ into value $\mathbf{V}=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$, where $\mathbf{W}_{p}$ and $\mathbf{W}_{d}$ respectively denote 1×1 point-wise convolution and 3×3 depth-wise convolution. We employ self-attention and get the output of PA-SA $\mathbf{\hat{F}}$:
$\hat{\mathbf{F}}=\mathbf{F}+softmax\left(\mathbf{Q}\mathbf{K}^{T}/\alpha\right)\cdot\mathbf{V},$ (13)
@@
where $\alpha$ is a learnable parameter. Consequently, PA-SA introduces physical guidance to fully exploit physics knowledge at the feature level and use self-attention mechanism to implicitly model the features of transmission map and background scattered light, which can help the diffusion model restore missing details and correct color distortion.
@@
Finaly, we design a multi-scale dynamic feed-forward network (MS-FFN) for local feature aggregation. In order to expand the receptive field, we employ multi-scale kernel depth-wise convolutions. MS-FFN adopts GELU to ensure the flexibility of feature aggregation. Thus, the multi-scale feature $\mathbf{F^{3}}$ and $\mathbf{F^{5}}$ of MS-FFN can be expressed as:
$\mathbf{F^{3}}$ $=\mathbf{W}^{3}_{d}\mathbf{W}_{p}(GELU(\mathbf{W}^{3}_{d}\mathbf{W}_{p}(Norm(\hat{\mathbf{F}})))),$ (14)
$\mathbf{F^{5}}$ $=\mathbf{W}^{5}_{d}\mathbf{W}_{p}(GELU(\mathbf{W}^{5}_{d}\mathbf{W}_{p}(Norm(\hat{\mathbf{F}})))),$
The output feature $F_{out}$ of MS-FFN can be expressed as:
$\mathbf{F_{out}}=Conv(Concat(\mathbf{F^{3}},\mathbf{F^{5}}))+\hat{\mathbf{F}}.$ (15)
@@
Table 1: Quantitative comparison on the UIEBD and LSUI datasets. The best results are highlighted in bold and the second best results are underlined.
@@
Fig. 2: Qualitative comparison with other SOTA methods on the UIEBD and LSUI datasets.
4 Experiments
4.1 Experimental Settings
Implementation details. Our network, implemented using PyTorch 1.7, underwent training and testing on an NVIDIA GeForce RTX 3090 GPU. We employed the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate was established at 0.0001. During the training stage, to balance the batch size and image size, the batch and image size are set to 8 and 128 × 128, respectively. The number of training iterations reached one million. The pixel values of the image are normalized to [-1,1]. The time step of the diffusion model is set to 2000. In the testing stage, following the setting in [7], the input size of the image is 256 × 256. By using the skip sampling strategy, and the sampling times are set to 10 times to balance the performance and runtime.
Datasets. We utilize the real-world UIEBD dataset [8] and the LSUI dataset [9] for training and evaluating our model. The UIEBD dataset comprises 890 underwater images with corresponding labels. Out of these, 700 images are allocated for training, and the remaining 190 are designated for testing. LSUI dataset contains 5004 underwater images and their corresponding high-quality images. Compared with UIEBD, LSUI contains diverse underwater scenes, object categories and deep-sea and cave images. In the paper, The LSUI dataset is randomly partitioned into 4500 images for training and 504 images for testing.
Evaluation Metrics. We employ standard full-reference image quality assessment metrics, including PSNR and SSIM [32], for quantitative comparisons at both pixel and structural levels. Higher PSNR and SSIM values indicate superior image quality. Additionally, LPIPS [33] and FID [34] are utilized to evaluate perceptual performance. Lower LPIPS and FID scores signify a more effective UIE approach.
4.2 Results and Comparisons
Table 1 shows the quantitative results compared with different SOTA methods on the UIEBD and LSUI datasets, including with UIECˆ2-Net [35], Water-Net [8], UIEWD [36], UWCNN [37], U-color [23], U-shape [9], and DM-water [7]. We mainly use PSNR, SSIM, LPIPS and FID as our quantitative indices for UIEBD and LSUI datasets. The results in Table 1 show that our algorithm outperforms state-of-the art methods obviously, especially in terms of perceptual metrics. To better validate the superiority of our methods, Fig. 2 shows the visual comparison results with other methods. The six examples are randomly selected on the UIEBD and LSUI datasets. Our methods consistently generate natural and better visual results, strongly proving that PA-Diff has good generalization performance for real-world underwater images.
4.3 Ablation Study
To evaluate the impact of each strategy on the diffusion model, we conduct an ablation study. Table 2 shows the ablation results on the LSUI dataset. Model A means that we just use the normal transformer structure. Compared with model A, model B, C and D achieve better results, suggesting that the knowledge of physics can help diffusion model to better restore image. Model B outperforms C, suggesting that the prior information of the transmission map is more important than the background light. Model E achieves the best performance, proving that our all contribution is effective for UIE tasks. Model E underperforms model F. The possible reason for this is that there is still a gap between the generated physics information (the transmission map and the background light) and the ideal true value. Therefore, how to generate more accurate a prior knowledge of physics remains a problem we need to research in the future.
Table 2: Ablation study on LSUI dataset. $T$ and $B$ refer to transmission map and background light prior information, respectively. $MS$ means multi-scale operation in our designed MS-FFN. $GAP$ denotes that we use global average pooling for the extraction of physical prior information from both the generated transmission map and background light.
@@ 4.3 Ablation Study
A $\times$ $\times$ $\times$ $\times$ 28.35 0.8943
B $\checkmark$ $\times$ $\times$ $\checkmark$ 24.71 0.9193
C $\times$ $\checkmark$ $\times$ $\checkmark$ 25.84 0.9077
D $\checkmark$ $\checkmark$ $\times$ $\checkmark$ 23.06 0.9298
E $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 25.44 0.9177
F $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 22.15 0.9354
5 Conclusion

### v2

@@ 3.2. Physics Prior Generation
(9) $\hat{I}=GT\cdot\mathcal{T}(I)+(1-\mathcal{T}(I))\mathcal{B}(GBB(I)),$
@@
We constrain the PPG branch by supervised manner. The physics reconstruction loss of PPG branch can be expressed as:
(10) $\mathcal{L}_{p}=||\hat{I}-I||_{1}.$
To enhance the perceptual similarity of the reconstructed images, we introduce the perceptual Loss $\mathcal{L}_{per}$ :
(11) $\mathcal{L}_{per}=||\phi(\hat{I})-\phi(I)||_{2},$
where $\phi$ is a pre-trained VGG network.
Finally, the overall loss $\mathcal{L}_{ppg}$ of PPG branch can be expressed as:
(12) $\mathcal{L}_{ppg}=\lambda_{1}\mathcal{L}_{p}+\lambda_{2}\mathcal{L}_{per}.$
Figure 3. The detailed structure of the proposed physics perception unit (PPU).
3.3. Implicit Neural Reconstruction
Inspired by (Quan et al., 2023; Chen et al., 2024; Yang et al., 2023a) based on implicit neural representations that can encode an image as a continuous function and learn a more robust feature representation, we further introduce Implicit Neural Representations (INRs) to learn robust degradation representations from diverse underwater image, aiming to reduce the difficulty of restoration for diffusion models. INRs trains a multi-layer perceptron (MLP) by learning the following mapping :
(13) $F_{\theta}:\mathbb{R}^{2}\to\mathbb{R}^{3},$
where the dimensions of input refer to the (x, y) spatial coordinates, while the dimensions of output refer to the (R, G, B) color channels. Specifically, following the setting of (Chen et al., 2024; Yang et al., 2023a), underwater image $I$ is transformed to a feature map $\mathbf{E}\in\mathbb{R}^{H\times W\times C}$ with a image resolution of H × W. Meanwhile, the position of each pixel is recorded in a relative coordinate set $\mathbf{X}\in\mathbb{R}^{H\times W\times 2}$, where the value ’2’ means horizontal and vertical coordinates. As shown in Figure 1, we fuse $X$ and $E$, and use $F_{\theta}$ to output image $\boldsymbol{I}_{{\mathrm{inr}}}$, which is expressed as:
(14) $\boldsymbol{I}_{{\mathrm{inr}}}[x,y]=\boldsymbol{F}_{\theta}(\mathbf{E}[x,y],\mathbf{X}[x,y]),$
where $[x,y]$ is the location of a pixel and $\boldsymbol{I}_{{\mathrm{inr}}}[x,y]$ is the generated RGB value. Due to the elimination of high-frequency components during rendering (Yang et al., 2023a), we adopt periodic spatial encoding (Lee and Jin, 2022) to project the pixel coordinates X into a higher dimensional space $\mathbb{R}^{2L}$ for better recovering high-frequency details. The encoding procedure is formulated as:
(15) $\mathbf{X}^{\prime}=\gamma(\mathbf{X}),$
(16) $\gamma(c)=(\cdot\cdot\cdot,\sin(2^{i}c),\cos(2^{i}c),\cdot\cdot\cdot),$
where $\gamma()$ represents a spatial encoding function. $i$ values from 00 to $L-1$, $c\in\mathbb{R}^{2L}$ is the coordinate value of $X$, and $L$ is a hyperparameter that determines the dimension value. Given a GT of underwater image $GT$, The reconstruction loss of INR branch can be expressed as:
(17) $\mathcal{L}_{INR}=||\boldsymbol{I}_{\mathrm{inr}}-GT||_{1}.$
Finally, we fuse $I$ and $\boldsymbol{I}_{\mathrm{inr}}$ to obtain the condition $x_{c}$ of diffusion model, which can be expressed as:
(18) $x_{c}=\boldsymbol{I}_{\mathrm{inr}}+I.$
3.4. Physics-aware Diffusion Transformer
@@
Physics-aware diffusion transformer (PDT) branch aims to exploit the extracted physics prior information from PPG branch for better guiding underwater image restoration. Fig. 1 shows the overall of PDT branch, which is a U-shaped structure with physics-aware diffusion transformer (PDT) blocks and cross attention module (CAM).
@@
Given the output of the INR branch as condition image $x_{c}\in{R}^{H\times W\times 3}$ and the background light $B^{c}\in{R}^{H\times W\times 3}$ as a physical prior condition, we first obtain the noise sample $x_{t}\in{R}^{H\times W\times 3}$ at time step $t$ according to the forward process, and concatenate $x_{t}$, $B^{c}$ and $x_{c}$ at channel dimension to obtain the input of PDT branch. Then, we obtain the embedding features $\mathbf{F}\in{R}^{H\times W\times C}$ through convolution, where C means the number of channel. $\mathbf{F}$ is encoded and decoded by PDT block, which consists of physics-aware self-attention (PA-SA), physics perception unit and gated multi-scale feed-forward network (GM-FFN). Note that, we introduce a physics perception unit (PPU) to further exploit the beneficial feature-level physics information.
Cross Attention Module. CAM aims to obtain a physics-aware interaction embedding, using which guides the diffusion process. We denote $T^{c}$ and $\mathbf{F}$ as the input of CAM. We use different linear projections to construct $Q$ and $K$ in CAM:
(19) $Q=Conv_{1\times 1}(T^{c}),$
(20) $K=Conv_{1\times 1}(\mathbf{F}).$
Similarly, $V$ of the transmission map $T^{c}$ can be obtained:
(21) $V=Conv_{1\times 1}(T^{c}).$
The output feature map of CAM $T_{out}$ can then be obtained from the formula:
(22) $T_{out}=Softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V.$
Physics-aware Self-attention. As is shown in Figure 2 (a), we first embed the time embedding $\mathbf{t}$ into the input features $\mathbf{F}$ in our PA-SA:
(23) $\tilde{\mathbf{F}}=Norm(\mathbf{F})+\mathbf{t},$
where $Norm$ means layer normalization. We first obtain query ($Q_{f}=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$), key ($K=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$), and value ($V=\mathbf{W}_{d}\mathbf{W}_{p}\tilde{\mathbf{F}}$) from the embedding input features $\mathbf{F}$. Meanwhile, we generate query ($Q_{t}=\mathbf{W}_{d}\mathbf{W}_{p}\mathbf{T_{out}}$) from the output feature map of CAM, where $\mathbf{W}_{p}$ and $\mathbf{W}_{d}$ respectively denote 1×1 point-wise convolution and 3×3 depth-wise convolution. We employ physics-aware query-driven attention strategy to obtain the final query $Q$ by fusing $Q_{t}$ and $Q_{f}$, which can be expressed as:
(24) $Q=Conv_{1\times 1}(concat(Q_{f},Q_{t})).$
We employ self-attention and get the output of PA-SA $\mathbf{\hat{F}}$:
(25) $\hat{\mathbf{F}}=\mathbf{F}+softmax\left(\mathbf{Q}\mathbf{K}^{T}/\alpha\right)\cdot\mathbf{V},$
@@
where $\alpha$ is a learnable parameter. Consequently, PA-SA introduces physical guidance to fully exploit physics knowledge at the feature level and use self-attention mechanism to implicitly model the features of transmission map, which can help the diffusion model restore missing details and correct color distortion.
Physics Perception Unit. Exploring feature-level physics information (Zheng et al., 2023; Dong and Pan, 2020) achieved success in image dehazing tasks. Our proposed physics perception unit (PPU) aims to further exploit the beneficial feature-level physics information for underwater image enhancement, and learn richer physical features. Our proposed PPU is based on the simplified atmospheric model Eq. (8), as shown in Fig. 3. We can reformulate the physics model to represent the clear image J, which can be obtained by:
(26) $J^{c}(\mathbf{x})=I^{c}(\mathbf{x})\frac{1}{T^{c}(\mathbf{x})}+B^{c}(1-\frac{1}{T^{c}(\mathbf{x})}).$
Let $k$ denote a feature extractor(the filter kernel in a deep network). Then extracting features via the feature extractor k, Eq. (25) can be reformulated as follows:
(27) $k\otimes J=k\otimes(I\odot\frac{1}{t})+k\otimes B(1-\frac{1}{t}),$
where $\otimes$ denotes the convolution operator and $\odot$ denotes the element-wise product operation. Using the matrix-vector form for algebraic operations, Eq. (26) can be express as:
(28) $\boldsymbol{KJ}=\boldsymbol{K}\boldsymbol{T}_{1}\boldsymbol{I}+\boldsymbol{K}\boldsymbol{T}_{2}\boldsymbol{B},$
where $\boldsymbol{K},\boldsymbol{J},\boldsymbol{I},\boldsymbol{B},\boldsymbol{T}_{1}$ and $\boldsymbol{T}_{2}$ respectively denote the matrix-vector forms of $k,J,I,B,\frac{1}{t}$ and $(1-\frac{1}{t})$. Note that the diagonal vector of the diagonal matrix $\boldsymbol{T}_{1}$ corresponds to the vectorized form of $\frac{1}{t}$. Assuming that $\boldsymbol{K}$ is an invertible matrix-vector, we can obtain
(29) $\boldsymbol{F}_{1}=\boldsymbol{K}\boldsymbol{T}_{1}\boldsymbol{K}^{-1},$
(30) $\boldsymbol{F}_{2}=\boldsymbol{K}\boldsymbol{T}_{2}\boldsymbol{K}^{-1}.$
And then, Eq. (27) can be express as:
(31) $\boldsymbol{KJ}$ $=(\boldsymbol{K}\boldsymbol{T}_{1}\boldsymbol{K}^{-1})\boldsymbol{K}\boldsymbol{I}+(\boldsymbol{K}\boldsymbol{T}_{2}\boldsymbol{K}^{-1})\boldsymbol{K}\boldsymbol{B}$
$=\boldsymbol{F}_{1}(\boldsymbol{K}\boldsymbol{I})+\boldsymbol{F}_{2}(\boldsymbol{K}\boldsymbol{B}).$
The Eq. (30) presents the relation between the clear image and the underwater image in a deep feature space. Due to the strong representation ability of deep neural networks, we can design a feature-level physics perception unit (PPU) and adopt deep networks to approximate the features corresponding to $\boldsymbol{F}_{1}$, $\boldsymbol{F}_{2}$ and $\boldsymbol{K}\boldsymbol{B}$. We denote $\tilde{\boldsymbol{B}}$ as an approximation of the features $\boldsymbol{K}\boldsymbol{B}$ that correspond to the background light, and $\tilde{\boldsymbol{t}_{1}}$ and $\tilde{\boldsymbol{t}_{2}}$ as approximations of $\boldsymbol{F}_{1}$ and $\boldsymbol{F}_{2}$, which are associated with the transmission map. Furthermore, $\boldsymbol{K}\boldsymbol{I}$ and $\boldsymbol{KJ}$ can be viewed as a latent underwater image $\tilde{\boldsymbol{I}}$ and its corresponding latent clear image $\tilde{\boldsymbol{J}}$, respectively. According to Eq. (30), the $\tilde{\boldsymbol{J}}$ can be express as:
(32) $\tilde{\boldsymbol{J}}=\tilde{\boldsymbol{t}_{1}}\odot\tilde{\boldsymbol{I}}+\tilde{\boldsymbol{t}_{2}}\odot\tilde{\boldsymbol{B}}.$
Since the background light is usually assumed to be homogeneous, we can use global average pooling (GAP) to obtain the main information for $\tilde{\boldsymbol{B}}$ in the feature space. $\tilde{\boldsymbol{B}}$ can be express:
(33) $\tilde{\boldsymbol{B}}=H(\sigma(\mathrm{Conv}(\mathrm{ReLU}(\mathrm{Conv}(\mathrm{GAP}(\tilde{\boldsymbol{I}}))))))))),$
where $\sigma$ is the Sigmoid function, H denotes a replication operation.
As the transmission map is non-homogeneous, we cannot apply GAP for the approximations of $\tilde{\boldsymbol{t}_{1}}$ and $\tilde{\boldsymbol{t}_{2}}$. We use different projection networks to obtain $\tilde{\boldsymbol{t}_{1}}$ and $\tilde{\boldsymbol{t}_{2}}$ :
(34) $\tilde{\boldsymbol{t}_{1}}=\sigma(\mathrm{Conv}(\mathrm{ReLU}(\mathrm{Conv}(\mathrm{Conv_{s}}(\tilde{\boldsymbol{I}})))))),$
(35) $\tilde{\boldsymbol{t}_{2}}=\sigma(\mathrm{Conv}(\mathrm{ReLU}(\mathrm{Conv}(\mathrm{Conv_{s}}(\tilde{\boldsymbol{I}})))))).$
where s represents the shared Conv layer. Given $\hat{\mathbf{F}}$ as the input of PPU, we use residual connection to obtain the final output of PPU, which can be expressed as :
(36) $\hat{\mathbf{F_{ppu}}}=PPU(\hat{\mathbf{F}})+\hat{\mathbf{F}}.$
@@
Gated Multi-scale Feed-Forward Network. Finaly, we design a GM-FFN for local feature aggregation. As is shown in Figure 2 (b), we employ GD-FFN-like structure (Zamir et al., 2022) to construct our feed-forward network. In order to expand the receptive field, we employ multi-scale kernel depth-wise convolutions. GM-FFN adopts GELU to ensure the flexibility of feature aggregation.
Figure 4. The visual results of the background scattered light $B^{c}$, the medium transmission map $T^{c}(x)$, and the implicit neural output $I_{inr}$.
@@
Table 1. Quantitative comparison on the UIEBD and LSUI datasets. The best results are highlighted in bold and the second best results are underlined.
@@
[…]
Figure 5. Qualitative comparison with other SOTA methods on the UIEBD and LSUI datasets.
Figure 6. More visual results from real-world scenarios.
Figure 7. Limitations. The output images refer to the reconstruction images of PPG branch.
4. Experiments
4.1. Experimental Settings
Implementation details. Our network, implemented using PyTorch 1.7, underwent training and testing on an NVIDIA GeForce RTX 3090 GPU. We employed the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate was established at 0.0001. During the training stage, to balance the batch size and image size, the batch and image size are set to 2 and 256 × 256, respectively. The number of training iterations reached one million. The pixel values of the image are normalized to [-1,1]. The time step of the diffusion model is set to 2000. In the testing stage, following the setting in (Tang et al., 2023), the input size of the image is 256 × 256. By using the skip sampling strategy, and the sampling times are set to 10 times to balance the performance and runtime.
Datasets. We utilize the real-world UIEBD dataset (Peng et al., 2023) and the LSUI dataset (Li et al., 2020) for training and evaluating our model. The UIEBD dataset comprises 890 underwater images with corresponding labels. Out of these, 700 images are allocated for training, and the remaining 190 are designated for testing. LSUI dataset contains 5004 underwater images and their corresponding high-quality images. Compared with UIEBD, LSUI contains diverse underwater scenes, object categories and deep-sea and cave images. In the paper, The LSUI dataset is randomly partitioned into 4500 images for training and 504 images for testing. In addition, to verify the generalization of PA-Diff, we use non-reference benchmarks U45 (Li et al., 2019), which contains 45 underwater images for testing.
Comparison methods. We conduct a comparative analysis between PA-Diff and seven state-of-the-art (SOTA) UIE methods, namely UIECˆ2-Net (Wang et al., 2021), Water-Net (Peng et al., 2023), UWCNN (Anwar et al., 2018), UIEWD (Ma and Oh, 2022), U-color (Li et al., 2021), U-shape (Li et al., 2020), and DM-water (Tang et al., 2023). To ensure a fair and rigorous comparison, we utilize the provided source codes from the respective authors and adhere strictly to the identical experimental settings across all evaluations.
Evaluation Metrics. We employ standard full-reference image quality assessment metrics, including PSNR and SSIM (Wang et al., 2004), for quantitative comparisons at both pixel and structural levels. Higher PSNR and SSIM values indicate superior image quality. Additionally, LPIPS (Zhang et al., 2018) and FID (Heusel et al., 2017) are utilized to evaluate perceptual performance. Lower LPIPS and FID scores signify a more effective UIE approach. For non-reference benchmarks U45, we introduce UIQM (Panetta et al., 2015) and UCIQE (Yang and Sowmya, 2015) to evaluate our method.
Table 2. Ablation study on LSUI dataset. $T^{c}$ and $B^{c}$ refer to transmission map and background light prior information in PPG branch, respectively. GBB means Gaussian blur module. INR represents our INR branch. PSE refers to the periodic spatial encoding (Lee and Jin, 2022), aiming to better recover high-frequency details. CAM and PPU means our designed cross attention module and physics perception unit , respectively. $GM-FFN$ means multi-scale operation in the feed-forward network.
[…]
A $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 28.35 0.8893
B $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 27.71 0.8974
C $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 25.92 0.9071
D $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 25.24 0.9112
E $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 24.50 0.9186
F $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ 24.92 0.9176
G $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ 23.66 0.9241
H $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ 22.85 0.9374
Ours $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 22.13 0.9378
4.2. Results and Comparisons
Table 1 shows the quantitative results compared with different SOTA methods on the UIEBD, LSUI and U45 datasets, including with UIECˆ2-Net (Wang et al., 2021), Water-Net (Peng et al., 2023), UIEWD (Ma and Oh, 2022), UWCNN (Anwar et al., 2018), U-color (Li et al., 2021), U-shape (Li et al., 2020), and DM-water (Tang et al., 2023). We mainly use PSNR, SSIM, LPIPS and FID as our quantitative indices for UIEBD and LSUI datasets, and UIQM and UCIQE for non-reference dataset U45. The results in Table 1 show that our algorithm outperforms state-of-the art methods obviously, especially in terms of perceptual metrics. To better validate the superiority of our methods, Fig. 5 shows the visual comparison results with other methods on the UIEBD and LSUI datasets. The some examples are randomly selected on the UIEBD and LSUI datasets. Furthermore, in order to more effectively validate the generalization performance of our method on real-world images, we present enhancement results from real scenes in Figure 6. Our methods consistently generate natural and better visual results, strongly proving that PA-Diff has good generalization performance for real-world underwater images.
4.3. Ablation Study
To evaluate the impact of each strategy on the diffusion model, we conduct a comprehensive ablation study. Table 2 shows the ablation results on the LSUI dataset. We use a full-reference assessment metric (SSIM) and perceptual assessment metric (FID) to evaluate the ablation study, respectively. Model A means that we just employ the most basic diffusion transformer structure. Firstly, we investigate the influence of the PPG branch on the diffusion process. Compared with model A, model B, C and D achieve better results, suggesting that the prior knowledge of physics can help diffusion model to better restore image. Particularly, Model B exhibits a remarkable advancement in terms of PSNR metric, suggesting that the transmission map $T^{c}$ plays a highly effective role in guiding the diffusion process. Model D outperforms C, suggesting that GBB is also useful to exploit more helpful the background light information for the diffusion process. Subsequently, we explore the impact of the INR branch on the performance of the diffusion model. Model E outperforms D, suggesting that INR branch can improve the performance and reduce the difficulty of restoration for diffusion models. Model F outperforms E, implying that the periodic spatial encoding can enhance the representation ability of INR branch. Subsequently, for models G and H, we respectively delve into the effects of our designed CAM and PPU on the model performance. Both models G and H yield improved outcomes, proving the significant role that CAM and PPU play in extracting more valuable feature-level physical information. Finally, Ours achieves the best performance, proving that our all contribution is effective for UIE tasks. Moreover, we also examine the applicability of PPU in other baseline models. Table 3 showcases the adaptability of PPU across different UIE methods, demonstrating its generalizability potential to exploit the beneficial feature-level physics information.
Limitations. As illustrated in Figure 7, there remains a discernible discrepancy between the reconstructed images from the PPG branch and the corresponding input images, indicating an existing gap between the physical information generated by our PPG branch and the ideal or ground-truth physical representation. Consequently, how to generate more accurate a prior knowledge of physics remains a problem we need to research in the future.
Table 3. Ablation study with the PPU on UIEBD.
Method Unet DW-water Ushape
w/o PPU(PSNR/SSIM) 17.41/0.7769 21.88/0.8194 21.25/0.8453
w PPU(PSNR/SSIM) 19.64/0.8157 21.23/0.8331 21.89/0.8570
5. Conclusion